OpenStudy (perl):

when is 2cos x + 1 >0 , < 0 AND = 0

OpenStudy (anonymous):

2cos x + 1 > 0 2cos x > -1 cos x > -1/2 x > arccos -1/2 2cos x + 1 < 0 2cos x < -1 cos x < -1/2 x < arccos -1/2 2cos x + 1 = 0 2cos x = -1 cos x = -1/2 x = arccos -1/2

OpenStudy (perl):

yes, all the solutions please

OpenStudy (anonymous):

arccos (-1/2) = 120 degrees, 240 degrees, (120+360) degrees, (240+360) degrees, (120+360+360) degrees, etc.

OpenStudy (perl):

increasing when 2pi*n < x < 2pi/3 + 2pi (n+1) , or 4pi/3 +2pi*n < x < 2pi + 2pi(n+1)

OpenStudy (perl):

solving when 2cos x + 1 > 0

OpenStudy (anonymous):

what is the entire question?

OpenStudy (perl):

A) Domain, B)Intercepts, C) symmetry, D)Asymptotes, E) Intervals, F) Intervals of Increase or Decrease, G) Concavity and Points of Inflection, H) Sketch the Curve y = sinx/(2 + cosx)

OpenStudy (anonymous):

what? that isn't a question.

OpenStudy (perl):

y = sinx/(2 + cosx find A) Domain, B)Intercepts, C) symmetry, D)Asymptotes, E) local max or min F) Intervals of Increase or Decrease, G) Concavity and Points of Inflection, H) Sketch the Curve

OpenStudy (anonymous):

ah, okay. Which one can't you do?

OpenStudy (perl):

all of em

OpenStudy (anonymous):

Well, I can't do some of them either. This is a different question from the first right?

OpenStudy (perl):

lets do intervals of increase and decrease

OpenStudy (perl):

here are the derivatives y = sinx/(2 + cosx) y ' = [ 2 cos (x) + 1 ] / ( cos x + 2)^2 y '' = [ 2 sin x ( cos x - 1) ] / ( cos (x) + 2) ^ 3

OpenStudy (anonymous):

http://www.youtube.com/watch?v=-W4d0qFzyQY

OpenStudy (perl):

increasing on (2pi*k, 2pi/3 + 2pi (k+1) union (4pi/3 +2pi*k, 2pi + 2pi(k+1) )

OpenStudy (perl):

do you agree?

OpenStudy (perl):

we know that y ' > 0 when 0 < x < 2pi/3 or 4pi/3 < x < 2pi, for x inside [ 0 , 2pi ]

OpenStudy (perl):

ok i messed up

OpenStudy (perl):

increasing on (2pi*k, 2pi/3 + 2pi*k) union (4pi/3 +2pi*k, 2pi + 2pi*k )

OpenStudy (perl):

we can make one nice solution if we do it is increasing on (-2pi/3 +2pi*k, 2pi/3 + 2pi*k ) and decreasing on (2pi/3 + 2pi*k , 4pi/3 +2pi*k )